# Linear spectroscopy of collective modes and the gap structure in two-dimensional superconductors

### v1

@@ Abstract
We consider optical response in multi-band, multi-layer two-dimensional superconductors. Within a simple model, we show that collective modes of the condensate, such as Leggett and clapping modes, can be detected in linear response. We show how trigonal warping of the superconducting order parameter can help facilitate detection of clapping modes. Taking rhombohedral trilayer graphene as an example, we consider several possible pairing mechanisms and show that all-electronic mechanisms may produce in-gap clapping modes. These modes, if present, should be detectable in the absorption of microwaves applied via gate electrodes; their detection would constitute strong evidence for unconventional pairing. Lastly, we show that absorption at frequencies above the superconducting gap $2|\Delta|$ also contains a wealth of information about the gap structure. Our results suggest that linear spectroscopy can be a powerful tool for the characterization of unconventional two-dimensional superconductors.
@@
In superconductors, a minimum of two collective modes arise: the Anderson-Bogoliubov-Goldstone (ABG) mode [1, 2, 3, 4], corresponding to phase fluctuations of the complex order parameter, and the Higgs mode [5, 6], corresponding to amplitude fluctuations. When the superconducting order parameter has more structure, its enlarged configuration space can produce a richer diversity of collective modes [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Measurements of the collective mode spectrum can therefore provide a valuable tool for diagnosing the underlying order. However, detection of exotic collective modes is often fraught with difficulties. Means for accessing nontrivial collective modes in linear response are therefore highly desirable. In this paper, we show how linear response to straightforward AC gating can reveal exotic collective modes in superconducting few-layer systems.
@@
As a prototypical example, consider the Leggett mode in multi-band superconductors [8]. One can think of this mode as a momentum-space analogue of the Josephson effect, consisting of relative number/phase fluctuations between the condensates formed from different bands. Assuming that intra-band interactions are attractive, when the Josephson coupling (i.e., inter-band pair swapping interaction) is not too strong, the Leggett mode can arise inside the superconducting gap, and it can therefore be underdamped. However, even when this mode is a well-defined excitation, crystal symmetries often render it invisible to linear electromagnetic response. The Leggett mode is therefore often difficult to access in experiments, requiring nonlinear response techniques such as Raman [23, 24, 25] or tunneling spectroscopy [26].
@@
Figure 1: (a) A two-layer superconductor under a large constant electric displacement field $D$, driven by a perturbing AC field $\delta D(t)$. (b) Single-electron dispersion in the toy model (black), showing the effect of a static perturbing field (red). (c) Schematic electron-electron scattering between Fermi sheets, indicating the relative phase ($\delta\theta$) and number ($\delta n$) fluctuations of the Leggett mode. (d) Absorption spectrum given $\nu_{1}=\nu_{2}$, $g_{11}^{(s)}=g_{22}^{(s)}$, and $\Delta_{1}=\Delta_{2}=\Delta$, for different values of interband coupling $g^{(s)}_{12}$ and $g^{(p)}=0$. (e, f) The clapping modes in a chiral $p$-wave superconductor, using the toy model with $g_{\alpha\alpha^{\prime}}^{(s)}=0>g^{(p)}$ and $|V|\gg\mu$. (e) Assuming a $p_{x}+ip_{y}$ ground state, the clapping modes at $\mathbf{q}\rightarrow 0$ correspond to fluctuations of the form $\Delta_{\mathbf{k}}(t)=\Delta_{0}e^{i\varphi_{\mathbf{k}}}+\left(a(t)+ib(t)\right)e^{-i\varphi_{\mathbf{k}}}$. Solid lines show a polar plot of the equilibrium order parameter, dashed lines show the distortion associated with each mode. (f) Real (dashed) and imaginary (solid) parts of the clapping mode contribution to the electronic compressibility for small but nonzero $\mathbf{q}$ in the circularly-symmetric model, Eq. (14); see text.
A simple gedankenexperiment illustrates the issue. Consider two identical superconducting layers (Fig. 1a) coupled by single-particle tunneling $J$. The tunnel coupling lifts the accidental degeneracy between the identical layers; suppose we are in a regime where electrons occupy both resulting bands. Direct driving of the Leggett mode requires forcing at least one of its conjugate variables, the relative density and relative phase between condensates. We could manipulate the relative phase by modulating the electric potential on each of the two layers with opposite sign, amounting to a displacement field $\mathbf{D}=D\hat{\mathbf{z}}$ out of the plane (Fig. 1a-c). However, before applying the field, the system is invariant under the mirror $M_{z}$, which exchanges the two layers (and maps $M_{z}\mathbf{D}=-\mathbf{D}$). This implies that the spectrum of the system is an even function of $D$, and for small fields there are thus no energy shifts at linear order; $E(D)=E(0)+O(D^{2})$. This picture also hints at what would solve the problem: perturbing around a static background field, $D=D_{0}+\delta D(t)$, in which case $D_{0}$ breaks $M_{z}$ and provides energy shifts linear in $\delta D$; $D^{2}=D_{0}^{2}+2D_{0}\delta D+(\delta D)^{2}$. We will show that this idea indeed works: the displacement field provides a useful and novel tool for enabling easy access to collective modes in few-layer superconductors. Moreover, the superconducting states in a wide variety of layered two-dimensional materials only arise in specific ranges of values of $D_{0}$ away from $D_{0}=0$, naturally giving access to their collective modes as described below.
@@
Besides the Leggett mode, another family of exotic collective modes arise in superconductors with spontaneously-broken time-reversal symmetry. Known as “generalized clapping modes” [9, 11, 12, 13], these amount to fluctuations in the uncondensed time-reversed partner of the condensed ground-state channel. For example, in a $p_{x}+ip_{y}$ chiral superconductor, clapping modes arise as fluctuations in the $p_{x}-ip_{y}$ channel, as shown in Fig. 1e. Direct observation of the clapping modes in linear electromagnetic response is typically not possible at zero wavevector, due to conservation of angular momentum. However, in superconducting rhombohedral trilayer graphene (RTG) [27], which may host a chiral $p$-wave order parameter, the coherence length is likely comparable to the device size. Therefore, finite-size effects may allow even a simple uniform gate to probe the clapping mode, as we will explore.
The rest of this paper is organized as follows. First, using a minimal toy model, we show how Leggett modes can be probed at $\mathbf{q}=0$ in systems with broken $M_{z}$ symmetry, and how clapping modes can be probed at small (but nonzero) $\mathbf{q}$ even when a high degree of rotational symmetry is present. We then consider a more realistic model of gated rhombohedral trilayer graphene (RTG). For the RTG model, we show how the observation of an in-gap collective mode can yield insight into the microscopic pairing interactions at play. We consider both conventional (phonon-mediated) and unconventional (electronic fluctuation-mediated) pairing mechanisms; only the unconventional mechanisms produce a chiral order parameter, and its associated clapping modes. We also show how the quasiparticle contribution to the linear response spectrum above $2|\Delta|$ can provide information on the superconducting gap structure.
While previous works have suggested that Leggett modes may aid in identifying the pairing mechanism in high-$T_{c}$ superconductors (see, e.g., Ref. [28]), our results show that in the case of 2D superconductors, AC gating can excite Leggett modes at linear order in the applied field, greatly simplifying possible experimental realizations. Additionally, other authors [22] have shown that clapping modes can modify the electronic compressibility at $\mathbf{q}\rightarrow 0$, but that this requires breaking of rotational symmetries. Working at finite $\mathbf{q}$, we focus on the experimentally-relevant case where (discrete) rotational symmetry is preserved, and show that the wavevector required for an appreciable signal corresponds to length scales comparable to device size in the case of RTG. Our results therefore point towards a new experimental tool for the difficult task of characterizing pairing in novel unconventional superconductors.
@@
In order to elucidate the basic ideas at work, we begin with a toy model which captures both the Leggett and clapping modes. The essential physics we are after is captured by the Hamiltonian $H=H_{0}+H_{\text{int}}$, with
@@
describing spin-1/2 fermions ($\sigma\in\{\uparrow,\downarrow\}$) in two layers ($l\in\{t,b\}$), each with bare dispersion $\xi_{\mathbf{k}0}=\epsilon_{\mathbf{k}0}-\mu$ relative to the Fermi level. We assume isotropic $\xi_{\mathbf{k}0}$ for simplicity. Single-particle tunneling ($J$) couples the layers, and the electric potential drops by $2V$ from top to bottom (we set the electron charge to $1$). Including interlayer tunneling, the band energies are $\xi_{\mathbf{k},\alpha\in\{1,2\}}=\xi_{\mathbf{k}0}\pm\sqrt{J^{2}+V^{2}}$. We assume a strong field, $|V|\gg|J|$, such that the band splitting is essentially linear in $|V|$. Our analysis can be applied to any pair of Fermi surfaces with arbitrary dispersions $\xi_{\mathbf{k}\alpha}$, so we leave these general for the most part.
@@
We base our analysis [29] on the imaginary-time path integral. To identify the collective modes, we perform a Hubbard-Stratonovich decoupling in the Cooper channel, and expand the effective action to quadratic order in fluctuations $\phi$ around the mean-field saddle point [30]. Analytically continuing the Gaussian effective action to real frequency ($i\Omega_{m}\rightarrow\Omega+i0^{+}$), the collective modes appear as poles of the bosonic propagator for pair fluctuations. To evaluate the linear response to applied electric fields, we also compute the couplings between the collective modes and such fields.
First we focus on the Leggett mode, by setting $g^{(p)}=0$ and $g^{(s)}_{\alpha\alpha^{\prime}}\neq 0$; afterwards we will instead set $g^{(s)}_{\alpha\alpha^{\prime}}=0$ and $g^{(p)}<0$ to focus on the clapping modes. The Leggett mode requires $\mu>\sqrt{J^{2}+V^{2}}$ such that both bands are occupied. Leggett’s original paper analyzed essentially this model—at $\mathbf{q}=0$ and with $g^{(p)}=0$—and found its relative-phase mode [8]. We assume $g_{\alpha\alpha}<0$. Taking $|\Delta_{1}|\leq|\Delta_{2}|$ without further loss of generality, we find [29] that an in-gap Leggett mode requires $0\leq|g^{(s)}_{12}|/(g^{(s)}_{11}g^{(s)}_{22}-|g^{(s)}_{12}|^{2})<\nu_{2}\arcsin(|\Delta_{1}/\Delta_{2}|)/\sqrt{1-(\Delta_{1}/\Delta_{2})^{2}}$, where $\nu_{\alpha}$ is the density of states per unit area per spin on the Fermi surface of band $\alpha$ in the normal state. (In the symmetric case where $|\Delta_{1}|=|\Delta_{2}|$, the second inequality is irrelevant).
We now show how an AC modulation of the displacement field can probe the Leggett mode. The drive field corresponds to a small time-dependent perturbation around the static background, $V\rightarrow V_{\rm DC}+\delta V(t)$. Keeping only terms up to linear order in $\delta V$ and working in the band basis, driving corresponds to a Hamiltonian term $H_{\delta V}=\sum_{\mathbf{k}\sigma}\delta V(t)(V_{\rm DC}/\sqrt{J^{2}+V_{\rm DC}^{2}})(c^{{\dagger}}_{\mathbf{k}1\sigma}c_{\mathbf{k}1\sigma}-c^{{\dagger}}_{\mathbf{k}2\sigma}c_{\mathbf{k}2\sigma})$. Integrating out the fermions in the presence of the drive yields a quadratic term $\sim\delta V^{2}$ and source terms $\sim\phi\,\delta V$ in the effective action, on top of the Gaussian action for pair fluctuations $\sim-\phi^{{\dagger}}\mathcal{D}^{-1}\phi$. Integrating out the fluctuations $\phi$ then yields the full quadratic free energy functional $\mathcal{F}[\delta V$], describing a capacitance (dependent on frequency $\Omega$):
@@
$\tilde{g}^{-1}=\frac{\left|g_{12}^{(s)}\right|}{g_{11}^{(s)}g_{22}^{(s)}-\left|g_{12}^{(s)}\right|^{2}}$ (6)
@@
$\gamma=\arcsin\left(\frac{\Omega+i0^{+}}{2\Delta}\right).$ (7)
@@
In the general (asymmetric) case, the key results are that at $\mathbf{q}\rightarrow 0$: i) The amplitude modes and ABG mode do not contribute to $\mathcal{C}(\Omega)$; and ii) The Leggett mode does contribute an in-gap pole to $\mathcal{C}(\Omega)$, provided that the background field $V_{\rm DC}\neq 0$. We make no attempt to explicitly compute the width of the pole, which is controlled by higher-order processes, and is model-dependent. However, we expect the resonance to be narrow, since no quasiparticle excitations are available to facilitate decay through low-order processes.
@@
To show how an AC electric field can probe the clapping modes, as in the Leggett case, we perturb around the static background, $V\rightarrow V_{\text{DC}}+\delta V(t)$. Since we projected out the upper band, the only effect of the modulation is to shift the energy of the lower band; when the displacement field-induced interlayer potential is very large compared with the interlayer hopping matrix element, the electrons only live on one layer, so they only feel the perturbing field through its local potential. Therefore, we calculate the clapping mode contribution to the electronic compressibility $\Pi^{00}$. Unlike the Leggett mode, the clapping modes carry angular momentum $2$ relative to the ground state; this means that with circular symmetry, their couplings to the scalar potential must vanish at $\mathbf{q}=0$. Hence, we calculate their compressibility contributions to leading nonvanishing order in $|\mathbf{q}|$.
@@
Here $\theta$ is the global phase (ABG) mode, $h$ the Higgs mode, and $a$ and $b$ the clapping modes. After integrating out the fermions, the long-wavelength action at quadratic order is
$S=\sum_{q}\Bigg{(}\sum_{\mu,\nu=0,x,y}\Pi^{\mu\nu}_{q}\mathcal{A}^{\mu}_{-q}\mathcal{A}^{\nu}_{q}\\ -\mathcal{D}^{-1}_{a,q}a_{-q}a_{q}-\mathcal{D}^{-1}_{b,q}b_{-q}b_{q}\\ +\sum_{\mu=0,x,y}\left[\Pi^{\mu a}_{q}\mathcal{A}^{\mu}_{-q}a_{q}+\Pi^{\mu b}_{q}\mathcal{A}^{\mu}_{-q}b_{q}\right]\Bigg{)},$ (12)
where $\mathcal{A}=(\mathcal{A}^{0},\bm{\mathcal{A}})=(A^{0}+\partial_{\tau}\theta,\mathbf{A}-\gradient\theta)$ is the gauge-invariant combination of the electromagnetic gauge potential $A$ and the spacetime gradient of the global phase $\theta$. Since we are only interested in the impact of the clapping modes on the compressibility, we drop all terms involving the spatial component $\bm{\mathcal{A}}$ for simplicity. Physically, we expect that this approximation should not drastically change the result, since the ABG mode associated with $\gradient\theta$ is either at a much lower energy (in a neutral superfluid) or much higher energy (after the Anderson-Higgs mechanism in a superconductor) than the clapping mode, for small but nonzero $\mathbf{q}$. We provide explicit expressions for the propagators $\mathcal{D}_{Xq}$ and couplings $\Pi^{0X}_{q}$ ($X\in\{a,b\}$) in the Supplemental Material [29]. Note that with circular symmetry, $\Pi^{0X}_{q}$ vanishes at $\mathbf{q}=0$ and first shows up at $O(|\mathbf{q}|^{2})$. The product $\Pi^{0X}_{q}\mathcal{A}^{0}_{-q}X_{q}$ must be scalar, so two factors of momentum are required to balance the angular momentum 2 of the clapping modes.
@@
$\delta\Pi^{00}_{q}\\ =\frac{\nu}{2}\left(\frac{\pi}{4}\xi_{\text{BCS}}|\mathbf{q}|\right)^{4}\csc^{2}\gamma\sec^{4}\gamma\frac{1-\gamma^{2}(\cot\gamma-\tan\gamma)^{2}}{\gamma(\cot\gamma-\tan\gamma)}.$ (14)
Here $\xi_{\text{BCS}}=v_{\text{F}}/(\pi\Delta_{0})$ is the BCS coherence length, $v_{\text{F}}=\sqrt{2\mu/m}$ is the Fermi velocity, and
$\gamma=\arcsin\left(\frac{\Omega+i0^{+}}{2\Delta_{0}}\right).$ (15)
This result neglects the dispersion of the clapping modes. Note the pole at $\cot\gamma=\tan\gamma$, corresponding to $\Omega=\pm\sqrt{2}\Delta_{0}$: the clapping modes are visible in the compressibility, and hence in the capacitance. In the vicinity of the positive-frequency pole,
@@
To recap, so far we have studied a toy model which captures both the Leggett and clapping modes. We showed that measurements of the AC capacitance (i.e., the response to a perturbing out-of-plane displacement field) can probe both of these classes of mode in 2D few-layer superconductors. Lastly, we pointed out that trigonal warping allows detection of the clapping modes at lower order in $|\mathbf{q}|$ by relaxing the absolute conservation of angular momentum to conservation mod 3, in keeping with $C_{3v}$ symmetry. We now move on to superconducting RTG as an illustrative case study.
We model the single-particle bandstructure of RTG using a $6\times 6$ tight-binding Hamiltonian in its continuum limit, written in the basis of two atoms per unit cell in each of the three layers [31, 32, 29]. We limit ourselves to the range of parameters (static perpendicular displacement field and doping level) where superconductivity was observed [27], focusing on the SC2 phase. In this regime, prior to the onset of superconductivity, RTG contains one annular Fermi sea in each valley (two in total), whose Fermi surfaces show substantial trigonal warping (see Fig. 2b). The specific pairing mechanism responsible for superconductivity in RTG is yet to be determined, so we consider conventional ($s$ wave, mediated by phonons) and unconventional (chiral $p$ wave, mediated by fluctuations of either charge density, as in the Kohn-Luttinger mechanism [33, 34, 35, 36], or intervalley coherence [37, 38]) scenarios [29].

### v2

@@ Abstract
We consider optical response in multi-band, multi-layer two-dimensional superconductors. Within a simple model, we show that linear response to AC gating can detect collective modes of the condensate, such as Leggett and clapping modes. We show how trigonal warping of the superconducting order parameter can help facilitate detection of clapping modes. Taking rhombohedral trilayer graphene as an example, we consider several possible pairing mechanisms and show that all-electronic mechanisms may produce in-gap clapping modes. These modes, if present, should be detectable in the absorption of microwaves applied via the gate electrodes, which are necessary to enable superconductivity in this and many other settings; their detection would constitute strong evidence for unconventional pairing. Lastly, we show that absorption at frequencies above the superconducting gap $2|\Delta|$ also contains a wealth of information about the gap structure. Our results suggest that linear spectroscopy can be a powerful tool for the characterization of unconventional two-dimensional superconductors.
@@
In superconductors, a minimum of two collective modes arise: the Anderson-Bogoliubov-Goldstone (ABG) mode [1, 2, 3, 4], corresponding to phase fluctuations of the complex order parameter, and the Higgs mode [5, 6], corresponding to amplitude fluctuations. When the superconducting order parameter has more structure, its enlarged configuration space can produce a richer diversity of collective modes [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. Measurements of the collective mode spectrum (e.g., by electromagnetic absorption) can therefore provide a valuable tool for diagnosing the nature of the order parameter, and even the underlying pairing mechanism [11, 12, 13, 33, 34, 35, 36, 37, 38, 30, 39, 40]. In this work, we show how to excite collective modes in gated two-dimensional superconductors using the gate electrodes themselves, providing a natural way to probe the nature of superconductivity in a wide range of emerging new materials.
@@
As a prototypical example, consider the Leggett mode a in multi-band superconductor [8]. One can think of this mode as a momentum-space analogue of the Josephson effect, consisting of relative number/phase fluctuations between the condensates formed from different bands. Assuming that intra-band interactions are attractive, when the Josephson coupling (i.e., inter-band pair swapping interaction) is not too strong, the Leggett mode can arise inside the superconducting gap, and it can therefore be underdamped. However, even when this mode is a well-defined excitation, crystal symmetries often render it invisible to linear electromagnetic response at $\mathbf{q}\rightarrow 0$ [41, 42]. As we show below, in a gated 2D superconductor, static gating can break $M_{z}$ mirror symmetry, permitting linear access to the Leggett mode at $\mathbf{q}\rightarrow 0$. The same gate electrodes can then be used to apply an AC probe field, thereby exciting the mode. This approach, which is uniquely well-suited to gated superconductors, represents an alternative to Raman [43, 34, 44, 21, 35, 28, 36, 45, 37, 38, 46, 47, 48, 39] and tunneling [33, 49] spectroscopy.
@@
Figure 1: (a) A two-layer superconductor under a large constant electric displacement field $D$, driven by a perturbing AC field $\delta D(t)$. (b) Single-electron dispersion in the toy model (black), showing the effect of a static perturbing field (red). (c) Schematic electron-electron scattering between Fermi sheets, indicating the relative phase ($\delta\theta$) and number ($\delta n$) fluctuations of the Leggett mode. (d) Absorption spectrum given $\nu_{1}=\nu_{2}$, $g_{11}^{(s)}=g_{22}^{(s)}$, and $\Delta_{1}=\Delta_{2}=\Delta$, for different values of interband coupling $g^{(s)}_{12}$ and $g^{(p)}=0$. (e, f) The clapping modes in a chiral $p$-wave superconductor, using the toy model with $g_{\alpha\alpha^{\prime}}^{(s)}=0>g^{(p)}$ and $|V|\gg\mu$. (e) Assuming a $p_{x}+ip_{y}$ ground state, the clapping modes at $\mathbf{q}\rightarrow 0$ correspond to fluctuations of the form $\Delta_{\mathbf{k}}(t)=\Delta_{0}e^{i\varphi_{\mathbf{k}}}+\left(a(t)+ib(t)\right)e^{-i\varphi_{\mathbf{k}}}$. Solid lines show a polar plot of the equilibrium order parameter, dashed lines show the distortion associated with each mode. (f) Real (dashed) and imaginary (solid) parts of the clapping mode contribution to the electronic compressibility for small but nonzero $\mathbf{q}$ in the circularly-symmetric model, Eq. (14); see text. We give $\Omega$ a small imaginary part by hand to give the pole a finite width.
@@
Besides the Leggett mode, another family of exotic collective modes arise in superconductors with spontaneously broken time-reversal symmetry. Known as “generalized clapping modes” [9, 10, 11, 13, 12, 14, 15, 16, 17, 30], these amount to fluctuations in the uncondensed time-reversed partner of the condensed ground-state channel. For example, in a $p_{x}+ip_{y}$ chiral superconductor, clapping modes arise as fluctuations in the $p_{x}-ip_{y}$ channel, as shown in Fig. 1e. Linear coupling of light to clapping modes at $\mathbf{q}\rightarrow 0$ generically requires breaking of rotational symmetry due to the non-vanishing angular momentum of these modes. Direct observation of clapping modes in linear electromagnetic response is therefore typically not possible at $\mathbf{q}\rightarrow 0$ in 2D materials with $C_{n}$ rotational symmetry, where $n>2$. Working at finite $\mathbf{q}$, we focus on the experimentally-relevant case where (discrete) rotational symmetry is preserved, and show that the wavevector required for an appreciable signal corresponds to length scales comparable to device size in the case of superconducting rhombohedral trilayer graphene (RTG), which may host a chiral $p$-wave order parameter [50]. Our results therefore point towards a particularly convenient method for characterizing pairing in gated 2D superconductors.
@@
In order to elucidate the basic ideas at work, we begin with a toy model which captures both the Leggett and clapping modes. The Hamiltonian is $H=H_{0}+H_{\text{int}}$, with
@@
describing spin-1/2 fermions ($\sigma\in\{\uparrow,\downarrow\}$) in two layers ($l\in\{t,b\}$), each with bare dispersion $\xi_{\mathbf{k}0}=\epsilon_{\mathbf{k}0}-\mu$ relative to the Fermi level. We assume isotropic $\xi_{\mathbf{k}0}$ for simplicity. Single-particle tunneling ($J$) couples the layers, and the electric potential drops by $2V$ from top to bottom (we set the electron charge to $1$). Including interlayer tunneling, the band energies are $\xi_{\mathbf{k},\alpha\in\{1,2\}}=\xi_{\mathbf{k}0}\pm\sqrt{J^{2}+V^{2}}$. Notice that starting from $V=0$, perturbing with a small $V$ produces energy shifts at second order in $V$. This reflects a more general principle: barring accidental degeneracies, any system which respects the mirror symmetry $M_{z}$ (which here exchanges the two layers) must have a spectrum which is even in the $M_{z}$-breaking out-of-plane displacement field, preventing linear access to the Leggett mode. In Fig. 1a, we assume a strong field ($|V|\gg|J|$), strongly breaking $M_{z}$, and rendering the band splitting essentially linear in $|V|$. This assumption is not essential to our calculations; any $V\neq 0$ will enable linear response to a subsequent perturbation $\delta V$.
@@
We base our analysis [51] on the imaginary-time path integral. To identify the collective modes, we perform a Hubbard-Stratonovich decoupling in the Cooper channel, and expand the effective action to quadratic order in fluctuations $\phi$ around the mean-field saddle point [52, 53, 54, 25, 28, 30]. Analytically continuing the Gaussian effective action to real frequency ($i\Omega_{m}\rightarrow\Omega+i0^{+}$), the collective modes appear as poles of the bosonic propagator for pair fluctuations. To evaluate the linear response to applied electric fields, we also compute the couplings between the collective modes and such fields.
First we focus on the Leggett mode, by setting $g^{(p)}=0$ and $g^{(s)}_{\alpha\alpha^{\prime}}\neq 0$. The Leggett mode requires $\mu>\sqrt{J^{2}+V^{2}}$ such that both bands are occupied. Leggett’s original paper analyzed essentially this model—at $\mathbf{q}=0$ and with $g^{(p)}=0$—and found its relative-phase mode [8]. We assume $g_{\alpha\alpha}<0$. Taking $|\Delta_{1}|\leq|\Delta_{2}|$ without further loss of generality, we find [51] that an in-gap Leggett mode requires $0\leq|g^{(s)}_{12}|/(g^{(s)}_{11}g^{(s)}_{22}-|g^{(s)}_{12}|^{2})<\nu_{2}\arcsin(|\Delta_{1}/\Delta_{2}|)/\sqrt{1-(\Delta_{1}/\Delta_{2})^{2}}$, where $\nu_{\alpha}$ is the density of states per unit area per spin on the Fermi surface of band $\alpha$ in the normal state. (When $|\Delta_{1}|=|\Delta_{2}|$, the second inequality is irrelevant). Fermi-liquid interactions can modify this condition [45]; see Sec. II of the Supplemental Material [51] for details.
We now show how an AC modulation of the displacement field can probe the Leggett mode. The drive field corresponds to a small time-dependent perturbation around the static background, $V\rightarrow V_{\rm DC}+\delta V(t)$. The physical idea is that in the presence of the static $V_{\text{DC}}\neq 0$ (which breaks $M_{z}$), the perturbation $\delta V$ produces linear energy shifts in opposite directions for the two bands, as illustrated in Fig. 1a, directly driving a phase difference between the two condensates and hence exciting the Leggett mode. Keeping only terms up to linear order in $\delta V$ and working in the band basis, driving corresponds to a Hamiltonian term $H_{\delta V}=\sum_{\mathbf{k}\sigma}\delta V(t)(V_{\rm DC}/\sqrt{J^{2}+V_{\rm DC}^{2}})(c^{{\dagger}}_{\mathbf{k}1\sigma}c_{\mathbf{k}1\sigma}-c^{{\dagger}}_{\mathbf{k}2\sigma}c_{\mathbf{k}2\sigma})$. Integrating out the fermions in the presence of the drive yields a quadratic term $\sim\delta V^{2}$ and source terms $\sim\phi\,\delta V$ in the effective action, on top of the Gaussian action for pair fluctuations $\sim-\phi^{{\dagger}}\mathcal{D}^{-1}\phi$. Integrating out the fluctuations $\phi$ then yields the full quadratic free energy functional $\mathcal{F}[\delta V$], describing a capacitance (dependent on frequency $\Omega$):
@@
$\tilde{g}^{-1}=\big{|}g_{12}^{(s)}\big{|}/\Big{(}g_{11}^{(s)}g_{22}^{(s)}-\big{|}g_{12}^{(s)}\big{|}^{2}\Big{)}$ (6)
@@
$\sin\gamma=(\Omega+i0^{+})/(2|\Delta|).$ (7)
@@
In the general (asymmetric) case, the key results are that: i) The amplitude modes and ABG mode (at $\mathbf{q}\rightarrow 0$) do not contribute to $\mathcal{C}(\Omega)$; and ii) The Leggett mode does contribute an in-gap pole to $\mathcal{C}(\Omega)$, provided that the background field $V_{\rm DC}\neq 0$ (and that the mode is in-gap). We make no attempt to explicitly compute the width of the pole, which is controlled by higher-order processes, and is model-dependent. However, we expect the resonance to be narrow, since no quasiparticle excitations are available to facilitate decay through low-order processes.
@@
To show how an AC electric field can probe the clapping modes, as in the Leggett case, we perturb around the static background field, $V\rightarrow V_{\text{DC}}+\delta V(t)$. Since we projected out the upper band, the only effect of the modulation is to shift the energy of the lower band; when the displacement field-induced interlayer potential is very large compared with the interlayer hopping matrix element, the electrons only live on one layer, so they only feel the perturbing field through its local potential. Therefore, we calculate the clapping mode contribution to the electronic compressibility $\Pi^{00}$. Unlike the Leggett mode, the clapping modes carry angular momentum $2$ relative to the ground state; this means that with circular symmetry, their couplings to the scalar potential must vanish at $\mathbf{q}=0$. Hence, we calculate their compressibility contributions to leading nonvanishing order in $|\mathbf{q}|$.
@@
Here $\theta$ is the global phase (ABG) mode, $h$ the Higgs mode, and $a$ and $b$ the clapping modes. After integrating out the fermions, and neglecting the (massive) Higgs mode for brevity, the long-wavelength action at quadratic order is
$S=\sum_{q}\Bigg{(}-\mathcal{D}^{-1}_{a,q}a_{-q}a_{q}-\mathcal{D}^{-1}_{b,q}b_{-q}b_{q}\\ +\Pi^{\mu\nu}_{q}\mathcal{A}^{\mu}_{-q}\mathcal{A}^{\nu}_{q}+\Pi^{\mu a}_{q}\mathcal{A}^{\mu}_{-q}a_{q}+\Pi^{\mu b}_{q}\mathcal{A}^{\mu}_{-q}b_{q}\Bigg{)},$ (12)
where $\mathcal{A}=(\mathcal{A}^{0},\bm{\mathcal{A}})=(A^{0}+\partial_{\tau}\theta,\mathbf{A}-\gradient\theta)$ is the gauge-invariant combination of the electromagnetic gauge potential $A$ and the spacetime gradient of the global phase $\theta$, and summation over $\mu,\nu\in\{0,x,y\}$ is implied. Since we are only interested in the impact of the clapping modes on the compressibility, we drop all terms involving the spatial component $\bm{\mathcal{A}}$ for simplicity. Physically, we expect that this approximation should not drastically change the result, since the ABG mode associated with $\gradient\theta$ is either at a much lower energy (in a neutral superfluid) or much higher energy (after the Anderson-Higgs mechanism in a superconductor) than the clapping mode, for small but nonzero $\mathbf{q}$. We provide explicit expressions for the propagators $\mathcal{D}_{Xq}$ and couplings $\Pi^{0X}_{q}$ ($X\in\{a,b\}$) in the Supplemental Material [51]. Note that with circular symmetry, $\Pi^{0X}_{q}$ vanishes at $\mathbf{q}=0$ and first shows up at $O(|\mathbf{q}|^{2})$. The product $\Pi^{0X}_{q}\mathcal{A}^{0}_{-q}X_{q}$ must be scalar, so two factors of momentum are required to balance the angular momentum 2 of the clapping modes.
@@
$\delta\Pi^{00}_{q}=2\nu\left(\frac{\pi}{4}\xi_{\text{BCS}}|\mathbf{q}|\right)^{4}\frac{1-\gamma^{2}(\cot\gamma-\tan\gamma)^{2}}{\gamma\cos^{2}\gamma\sin 4\gamma}.$ (14)
Here $\xi_{\text{BCS}}=v_{\text{F}}/(\pi\Delta_{0})$ is the BCS coherence length, and $v_{\text{F}}=\sqrt{2\mu/m}$ is the Fermi velocity. This result neglects the dispersion of the clapping modes. Note the pole at $\gamma=\pi/4$, corresponding to $\Omega=\pm\sqrt{2}\Delta_{0}$: the clapping modes are visible in the compressibility, and hence in the capacitance. In the vicinity of the positive-frequency pole,
@@
To go beyond our simple toy model, we consider superconducting RTG as an illustrative case study. We model the single-particle bandstructure of RTG using a $6\times 6$ tight-binding Hamiltonian in its continuum limit, written in the basis of two atoms per unit cell in each of the three layers [55, 56, 51]. We limit ourselves to the range of parameters (static perpendicular displacement field and doping level) where superconductivity was observed [50], focusing on the SC2 phase. In this regime, prior to the onset of superconductivity, RTG contains one annular Fermi sea in each valley (two in total), whose Fermi surfaces show substantial trigonal warping (see Fig. 2b). The specific pairing mechanism responsible for superconductivity in RTG is yet to be determined, so we consider conventional ($s$ wave, mediated by phonons) and unconventional (chiral $p$ wave, mediated by fluctuations of either charge density, as in the Kohn-Luttinger mechanism [57, 58, 59, 60], or intervalley coherence [61, 62]) scenarios [51]. We construct a general formalism for calculating superconducting collective modes and their response functions in the Supplemental Material [51].